OpenStudy (micheal1234):

Identity 16. Solve for the variable: 3n – 5 = 7n + 11 17. Solve for the variable: -6a + 3 = -3(2a – 1) 18. What is the solution of -(-5 – 6x) = 4(5x + 3)? A. -2 B. -1/2 C. ½ d. 2

OpenStudy (michaelac959):

i got this

OpenStudy (photon336):

you can add 5 to both sides 3n-5 = 7n+11 Subtract 7n from both sides 3n = 7n+16 -7n divide both sides by four Try to attempt the second one on your own and post what you got. -4n = 16 -4 -4 n = -4 CHECK 3n = 7n+16 3(-4) = -12 7(-4) = -28+16 = -12 -12 = -12 that's how we know this answer is correct. so n = -4

OpenStudy (micheal1234):

thanks

OpenStudy (michaelac959):

wow im slow

OpenStudy (micheal1234):

only need 17 and 18